OpenStudy (anonymous):

If X = (4,0) what does X' equal if Dz,2?

OpenStudy (anonymous):

Dz ?

OpenStudy (anonymous):

I don't get it either. That is just what the problem said.

OpenStudy (anonymous):

Hmm maybe: X = x(x -4) X = x^2 - 4x X' = 2x - 4 X' = 2(x-2) But i'm not sure what Dz is ment to mean. Is the above the sort of stuff you've been doing?

OpenStudy (anonymous):

Yes it is, that is why I don't understand why the Dz suddenly came up. Thank you though!